# TLOB: A Novel Transformer Model with Dual Attention for Stock Price Trend Prediction with Limit Order Book Data

### v2

@@ 7.2. Tesla and Intel results
Table 4. F1-score for Intel on four horizons. Bold values show the best scores.
@@
In Table 3 we show the results for Tesla and in Table 4 for Intel. For each stock, we trained a different model. In the Appendix (B) we report also the precision and recall curves for a horizon equal to 100. For INTC, they exhibit excellent precision at low recall values, indicating their ability to accurately identify the most confident positive instances. MLPLOB outperforms every model on the first two horizons (10, 20), while on the longer horizons (50, 100) TLOB outperforms every model. This is expected since Transformers excels at long-range dependencies. Notably, the difference in performance between MLPLOB and TLOB for the shorter horizons is minimal ($\approx 0.5$), while on the longer horizons is significant ($\approx 7$). As expected the longer the horizon the more difficult to forecast. In general, the performances are much lower with respect to FI-2010. We conjecture that this is due to the fact that FI-2010 is characterized by a lower level of complexity with respect to NASDAQ stocks. This derives from the fact that it is composed of Finnish stocks, which are less liquid and efficient than NASDAQ stocks such as Intel and Tesla. Additionally, the data dates back to 2010. Indeed, as will be demonstrated in the subsequent experiment, the prediction difficulty augments as time goes by. The results of our experiment are supported by several works on the topic (Prata et al., 2024; Zhang et al., 2019; Sirignano, 2019). All the models are trained until convergence. Notably, both TLOB and MLPLOB achieve convergence in less than half the epochs required by BiNCTABL and DeepLOB.
7.3. Are stocks harder to forecast than in the past?
Table 5. F1-score for Intel on two different periods, from 2012 and 2015. The horizon is set to 50.
@@ 7.3. Are stocks harder to forecast than in the past?
This experiment examines the challenges associated with market prediction over time and the self-destruction of predictable patterns in financial markets. Empirical evidence consistently demonstrates that forecasting models effective in certain periods become obsolete over time. Several studies indicate that previously observed predictability patterns disappeared after becoming widely known. Dimson and Marsh (Dimson and Marsh, 1999) found this for the UK small-cap premium, while Bossaert and Hillion (Bossaerts and Hillion, 1999) noted a decline in international stock return predictability around 1990. Aiolfi and Favero (Aiolfi and Favero, 2005) reported similar findings for US stocks in the 1990s. The market is increasingly efficient and difficult to predict as time goes by. We extend this investigation to our best-performing model TLOB. Specifically, we tested on a day of Intel from 2012/06/21777we remark that in a single day of Intel, there are hundreds of thousands of orders making the experiment statistically significant. Furthermore, the trading day was extracted from the LOBSTER public sample files available at https://lobsterdata.com/info/DataSamples.php and it was the only day available, eliminating the possibility of cherry picking. and confronted the difference in performance with 2015/01/30. We report the performance in Table 5. As expected the performance from 2012 is better than that from 2015. We confirm the hypothesis and the empirical evidence from other works.
7.4. Alternative Threshold Definition Using Average Spread
Table 6. F1-score on Tesla with $\theta$ set to the average spread.
@@ 7.4. Alternative Threshold Definition Using Average Spread
Based on the fact that predictability has to be considered in relation to the transaction costs, we explore an alternative approach to define the trend classification parameter $\theta$, setting it equal to the average spread as a percentage of the mid-price, reflecting the primary transaction cost. This methodology could only be applied to Tesla data, as Intel’s higher trading volume (approximately 10 times greater in January 2015) and lower volatility relative to traded shares would result in 99.99% of trends classified as stationary. We set the horizons to 50, 100, and 200 because with shorter horizons 99% of the mid-price movements would be classified as stationary. In Table 6 we report the results. In general, performances show a deterioration, which is probably caused by the classes’ unbalance. This experiment highlights the necessity for further refinements in trend definition and method complexity when targeting profitability in practical applications.
7.5. Ablation Study
To evaluate the contribution of each attention mechanism within the TLOB architecture, we performed an ablation study on the FI-2010 dataset. Specifically, we compared the performance of the complete TLOB model against two ablated versions: one without spatial attention (TLOB w/o SA) and another without temporal attention (TLOB w/o TA). To avoid inconsistency, we maintain the total number of layers fixed [^8] TLOB has 4 temporal attention layers and 4 spatial attention layers, TLOB w/o SA has 8 temporal attention layers and TLOB w/o TA has 8 spatial attention layers. The F1-scores for each model across four prediction horizons (h = 10, 20, 50, and 100) are presented in Table 7. The results demonstrate that the full TLOB model, incorporating both spatial and temporal attention mechanisms, consistently outperforms both ablated versions across all prediction horizons. The performance gain of the full TLOB model highlights the importance of capturing both spatial relationships between LOB features and temporal dependencies across LOB snapshots. This suggests that the dual-attention mechanism effectively learns complementary information, leading to improved predictive accuracy compared to models relying on only one type of attention.
Table 7. Ablation study results. F1-score on the FI-2010 dataset on four horizons. Bold values show the best scores.
@@ 7.5. Ablation Study
8. Conclusion
We proposed two new deep learning models MLPLOB: A simplified yet effective MLP-based architecture and TLOB, a Transformer-based approach, for the task of stock price trend prediction on Limit Order Book (LOB) data. Both models demonstrated superior performance compared to existing state-of-the-art approaches, with TLOB showing particular promise in handling high-frequency market data. NASDAQ stocks (Tesla, Intel) proved significantly more challenging to predict than Finnish stocks (FI-2010). Our research also showed that prediction accuracy decreases as the forecasting horizon increases, highlighting the inherent challenges of long-term prediction in financial markets.
Limitations: When considering practical implementation, we found that defining trend thresholds based on average spread (transaction costs) significantly impacts model evaluation and potential profitability. This finding underscores the critical gap between academic performance metrics and practical trading applicability.
Future works Looking ahead, several avenues for future research emerge. The investigation of scaling laws for financial deep learning models remains an open question, as does the development of more robust approaches to handling increased market efficiency and complexity. Additionally, the exploration of alternative trend definition methodologies that better align with practical trading constraints could prove fruitful.
Risks: Firstly, it is important to acknowledge that the proposed methodologies are not sufficiently mature for practical deployment in live trading environments. Furthermore, the application of deep learning models to stock price prediction and subsequent utilization in trading carries significant inherent risks. A primary concern is the limited explainability of such models. Furthermore, automated AI models, increasingly integrated into financial markets, present significant risks to financial stability due to their potential to amplify systemic vulnerabilities. These models, often operating with limited transparency, can trigger rapid and widespread market reactions, exacerbating volatility and potentially leading to cascading failures across the financial system.
@@ Appendix A Hyperparameters Search
To find the best hyperparameters, we employ a grid search exploring different values as shown in Table 8. Regarding the hyperparameters of DeepLOB and BiNCTABL, we used the one used in (Prata et al., 2024) after a large hyperparameters search. We remark that with higher sequence sizes than 128, the performances reach a plateau. For TLOB we searched also for the optimal number of heads, and we noted that there was not difference in between performance between 1, 2, 4 and so we fixed the number of heads to 1.
Table 8. The hyperparameter search spaces and best choices for each model.
@@
Optimizer {Adam (Diederik, 2014), Lion (Chen et al., 2024)} Adam Adam
Sequence size {64, 128, 256, 384, 512} 128 128
@@ Appendix B Additional Results
We report the precision and recall curves for FI-2010 (Fig. 4), INTC (Fig. 5) and TSLA (Fig. 6), for horizon 100. As shown, across the different datasets, TLOB and MLPLOB consistently achieve higher precision at all recall levels compared to the other models. TLOB and MLPLOB, for INTC, exhibit excellent precision at low recall values, indicating their ability to accurately identify the most confident positive instances. Specifically for FI-2010, they exhibit very high precision, also at high recall values.
Figure 5. Precision and Recall curve for INTC for horizon = 100.
Figure 6. Precision and Recall curve for TSLA for horizon = 100.

### v3

@@ 7.2 Tesla and Intel results
Table 5: F1-score for Intel on four horizons. Bold values show the best scores.
@@
In Table 4 we show the results for Tesla and in Table 5 for Intel. For each stock, we trained a different model. In the Supplementary Material, we also report the precision and recall curves for a horizon equal to 100. For INTC, they exhibit excellent precision at low recall values, indicating their ability to accurately identify the most confident positive instances. MLPLOB outperforms every model on the first two horizons (10, 20), while on the longer horizons (50, 100), TLOB outperforms every model. This is expected since Transformers excels at long-range dependencies. Notably, the difference in performance between MLPLOB and TLOB for the shorter horizons is minimal ($\approx 0.5$), while on the longer horizons it is significant ($\approx 7$). As expected, the longer the horizon, the more difficult to forecast. In general, the performances are much lower with respect to FI-2010. We conjecture that this is due to the fact that FI-2010 is characterized by a lower level of complexity with respect to NASDAQ stocks. This derives from the fact that it is composed of Finnish stocks, which are less liquid and efficient than NASDAQ stocks such as Intel and Tesla. Additionally, the data of FI-2010 dates back to 2010. Indeed, as will be demonstrated in the subsequent experiment, the prediction difficulty augments as time goes by. All the models are trained until convergence. Notably, both TLOB and MLPLOB achieve convergence in less than half the epochs required by BiNCTABL and DeepLOB.
7.3 BTC Results
In Table 6 we show the results for the BTC dataset. We remark that this dataset is from 2023, so it is the most recent one  [^9] FI-2010 is from 2010, while TSLA-INTC is from 2015. Unlike the other datasets, TLOB outperforms every model on every horizon. This consistent dominance on the most recent (2023) dataset, particularly with a volatile asset like Bitcoin, suggests TLOB’s architecture is highly effective at capturing contemporary market dynamics that may elude other models. Moreover, the widening performance gap at longer horizons (h=50, 100) with respect to MLPLOB hints at TLOB’s enhanced capacity to model the complex, longer-term temporal dependencies. Similarly to the other two datasets, the difference in performance between MLPLOB and TLOB for the shorter horizons (10, 20) is minimal ($\approx 0.5$), while on the longer horizons it is significant ($\approx 5$).
Table 6: F1-score for BTC on four horizons. Bold values show the best scores.
BTC F1-Score (%) $\uparrow$
Model h = 10 h = 20 h = 50 h = 100
DeepLOB 68.07 57.87 45.13 37.43
BiNCTABL 73.4 61.34 47.05 40.59
MLPLOB 74.6 61.02 42.74 36.97
TLOB 74.7 61.74 48.54 41.49
7.4 Are stocks harder to forecast than in the past?
Table 7: F1-score for Intel on two different periods, from 2012 and 2015. The horizon is set to 50.
@@ 7.4 Are stocks harder to forecast than in the past?
This experiment examines the challenges associated with market prediction over time and the self-destruction of predictable patterns in financial markets. Empirical evidence consistently demonstrates that forecasting models effective in certain periods become obsolete over time. Several studies indicate that previously observed predictability patterns disappeared after becoming widely known. Dimson and Marsh [16] found this for the UK small-cap premium, while Bossaert and Hillion [3] noted a decline in international stock return predictability around 1990. Aiolfi and Favero [1] reported similar findings for US stocks in the 1990s. The market is increasingly efficient and difficult to predict as time goes by. We extend this investigation to our best-performing model, TLOB. Specifically, we tested on a day of Intel from 2012/06/21101010we remark that in a single day of Intel, there are hundreds of thousands of order,s making the experiment statistically significant. Furthermore, the trading day was extracted from the LOBSTER public sample files available at https://lobsterdata.com/info/DataSamples.php, and it was the only day available, eliminating the possibility of cherry picking. and confronted the difference in performance with 2015/01/30. We report the performance in Table 7. As expected, the performance from 2012 is better than that from 2015. We confirm the hypothesis and the empirical evidence from other works.
7.5 Alternative Threshold Definition Using Average Spread
Table 8: F1-score on Tesla with $\theta$ set to the average spread.
@@ 7.5 Alternative Threshold Definition Using Average Spread
Based on the fact that predictability has to be considered in relation to the transaction costs, we explore an alternative approach to define the trend classification parameter $\theta$, setting it equal to the average spread as a percentage of the mid-price, reflecting the primary transaction cost. This methodology could only be applied to Tesla data, as Intel’s higher trading volume (approximately 10 times greater in January 2015) and lower volatility relative to traded shares would result in 99.99% of trends classified as stationary. We set the horizons to 50, 100, and 200 because with shorter horizons, 99% of the mid-price movements would be classified as stationary. In Table 8 we report the results. In general, performances show a deterioration, which is probably caused by the classes’ unbalance. This experiment highlights the necessity for further refinements in trend definition and method complexity when targeting profitability in practical applications.
7.6 Ablation Study
To evaluate the contribution of each attention mechanism within the TLOB architecture, we performed an ablation study on the FI-2010 dataset. Specifically, we compared the performance of the complete TLOB model against two ablated versions: one without spatial attention (TLOB w/o SA) and another without temporal attention (TLOB w/o TA). To avoid inconsistency, we maintain the total number of layers fixed [^11] TLOB has 4 temporal attention layers and 4 spatial attention layers, TLOB w/o SA has 8 temporal attention layers, and TLOB w/o TA has 8 spatial attention layers. The F1-scores for each model across four prediction horizons (h = 10, 20, 50, and 100) are presented in Table 9. The results demonstrate that the full TLOB model, incorporating both spatial and temporal attention mechanisms, consistently outperforms both ablated versions across all prediction horizons. The performance gain of the full TLOB model highlights the importance of capturing both spatial relationships between LOB features and temporal dependencies across LOB snapshots. This suggests that the dual-attention mechanism effectively learns complementary information, leading to improved predictive accuracy compared to models relying on only one type of attention.
Table 9: Ablation study results. F1-score on the FI-2010 dataset on four horizons. Bold values show the best scores.
@@ 7.6 Ablation Study
8 Conclusion
We proposed MLPLOB and TLOB, two deep-learning architectures for LOB-based price trend prediction. Both outperformed SoTA methods on FI-2010 (Finnish stocks), NASDAQ stocks (Tesla, Intel), and Bitcoin, especially for longer horizons. When considering practical implementation, we found that defining trend thresholds based on average spread (transaction costs) significantly impacts model performance and profitability. This finding underscores the critical gap between machine learning metrics and practical trading applicability.
Future works: Looking ahead, several avenues for future research emerge. The investigation of scaling laws for financial deep learning models remains an open question, as does the development of more robust approaches to handling increased market efficiency and complexity. Additionally, the exploration of alternative trend definition methodologies that better align with practical trading constraints could prove fruitful. Finally an extensive profitability analysis, based on backtesting or other more advanced market simulations methods [10, 11, 2], would be very interesting.
Limitations: Firstly, it is important to acknowledge that the proposed methodologies are not sufficiently mature for practical deployment in live trading environments. When considering practical implementation, we found that defining trend thresholds based on average spread (transaction costs) significantly impacts model evaluation and potential profitability. This finding underscores the critical gap between academic performance metrics and practical trading applicability.
Risks: Automated ML models, increasingly integrated into financial markets, present significant risks to financial stability due to their potential to amplify systemic vulnerabilities. These models can trigger rapid and widespread market reactions, exacerbating market volatility and potentially leading to cascading failures across the financial system.
@@ Appendix A Hyperparameters Search
To find the best hyperparameters, we employ a grid search exploring different values as shown in Table 10. Regarding the hyperparameters of DeepLOB and BiNCTABL, we used the ones used in [31] after a large hyperparameter search. We remark that with higher sequence sizes than 128, the performances reach a plateau. For TLOB, we also searched for the optimal number of heads, and we noted that there was not difference in between performance between 1, 2, 4, and so we fixed the number of heads to 1.
Table 10: The hyperparameter search spaces and best choices for each model.
@@
Optimizer {Adam [15], Lion [9]} Adam Adam
Sequence size {64, 128, 256, 384, 512} 128 384
@@ Appendix B Additional Results
We report the precision and recall curves for FI-2010 (Fig. LABEL:fig:pr_fi), INTC (Fig. 3), and TSLA (Fig. 4), for horizon 100. As shown, across the different datasets, TLOB and MLPLOB consistently achieve higher precision at all recall levels compared to the other models. TLOB and MLPLOB, for INTC, exhibit excellent precision at low recall values, indicating their ability to accurately identify the most confident positive instances. Specifically for FI-2010, they exhibit very high precision, also at high recall values.
Figure 3: Precision and Recall curve for INTC for horizon = 100.
Figure 4: Precision and Recall curve for TSLA for horizon = 100.